The following are the bundles that Akua can afford with her income:

Yoghurt

Bread

10

 

 


 

80

20

 

Fill the blanks in the table with the quantities that will exhaust Akua’s income.

 



1
Expert's answer
2021-05-09T14:27:39-0400

Yoghurt Bread

10 160

50 0

30 80

20 120


Total Income = 400 cedis.

With 400 cedis, Akua can purchase either 50 units of Yogurts or 200 Loaves of bread.

Price of Yogurt per unit = Total Income/ Maximum units of Yogurt consumed = 400/50 = 8 cedis

Price of one Loave of bread  = Total Income/ Maximum units of loaves of bread consumed = 400/200 = 2 cedis

Now, we can complete the table.

If Akua buys 10 units of Yogurt then she has exhausted 10* Price of Yogurt per unit = 10*8 = 80 cedis.

Remaining income for consumption of Loaves of bread = 400 - 80 = 320 cedis

Therefore, Akua can buy 320/2 = 160 loaves of bread maximum.

If Akua buys 0 units of Loaves then she has all the income available to spend on Yogurt.

Available income for consumption of Yogurt = 400 cedis

Therefore, Akua can buy 400/8 = 50 units of Yogurt, maximum.

If Akua buys 80 units of Loaves then she has exhausted 80* Price of loave of bread = 80*2 = 160 cedis.

Remaining income for consumption of Yogurt = 400 - 160 = 240 cedis

Therefore, Akua can buy 240/8 = 30 units of Yogurt maximum.

If Akua buys 20 units of Yogurt then she has exhausted 20* Price of Yogurt per unit = 20*8 = 160 cedis.

Remaining income for consumption of Loaves of bread = 400 - 160 = 240 cedis

Therefore, Akua can buy 240/2 = 120 loaves of bread maximum.